# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ Abstract
This paper investigates the detection and estimation of a single change in high-dimensional linear models. We derive minimax lower bounds for the detection boundary and the estimation rate, which uncover a phase transition governed the sparsity of the covariance-weighted differential parameter. This form of inherent sparsity captures a delicate interplay between the covariance structure of the regressors and the change in regression coefficients on the detectability of a change point. Complementing the lower bounds, we introduce two covariance scanning-based methods, McScan and QcSan, which achieve minimax optimal performance (up to possible logarithmic factors) in the sparse and the dense regimes, respectively. In particular, QcScan is the first method shown to achieve consistency in the dense regime and further, we devise a combined procedure which is adaptively minimax optimal across sparse and dense regimes without the knowledge of the sparsity. Computationally, covariance scanning-based methods avoid costly computation of Lasso-type estimators and attain worst-case computation complexity that is linear in the dimension and sample size. Additionally, we consider the post-detection estimation of the differential parameter and the refinement of the change point estimator. Simulation studies support the theoretical findings and demonstrate the computational and statistical efficiency of the proposed covariance scanning methods.
@@ 1 Introduction
The change point problem in linear models has received great attention over the past several decades, owing to its widespread applicability across diverse scientific and practical disciplines, including genomics, climatology, finance and economics. Numerous methods have been developed that achieve consistency in (multiple) change point detection, both in multivariate settings (e.g. Csörgo and Horváth,, 1997; Bai and Perron,, 1998) and high dimensional scenarios (e.g. Lee et al.,, 2016; Kaul et al.,, 2019; Wang et al., 2021b, ; Rinaldo et al.,, 2021; Cho and Owens,, 2024; Xu et al.,, 2024).
When faced with the high dimensionality, many existing methods rely on $\ell_{1}$-regularised estimators of local regression parameters for change point detection, thereby requiring the sparsity of regression parameters. Gao and Wang, (2022) propose CHARCOAL, a projection-based method that instead assumes sparsity on the differences of regression parameters. However, CHARCOAL is applicable only when the dimension is strictly smaller than the sample size. More recently, Cho et al., (2025) introduce McScan, a multiscale covariance scanning approach that achieves consistent change point detection with both statistical and computational efficiency. McScan does not require exact sparsity of the regression parameters or their differences and accommodates ultra-high dimensional settings. It operates by examining the $\ell_{\infty}$-norm of the contrasts in local sample covariances between regressors and responses, and is thus implicitly better suited to sparse regimes. Minimax lower bounds for the change point problem have been investigated, but only in the sparse setting when the $\ell_{0}$-norm of regression parameters (or their differences) is bounded by the square root of the dimension (see Rinaldo et al.,, 2021 and Cho et al.,, 2025).
Despite these advances, several questions remain to be answered for the change point problem in high dimension linear models. In particular, it is hitherto unknown whether it is possible to consistently detect the change point when the change is non-sparse. More fundamentally, it is yet to be understood which notion of sparsity — or more generally, which structural assumption — determines the intrinsic difficulty in the detection and estimation of change points. In this paper, we provide first answers to these questions under a Gaussian linear model with at most one change (AMOC): The observations $(Y_{t},\mathbf{x}_{t})\in\mathbb{R}^{1+p},\,t\in[n]:=\{1,\ldots,n\}$, are distributed as
@@
where $\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})$ with non-negative definite $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ and independently, $\varepsilon_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2})$ for all $t$. We denote the differential parameter by $\bm{\delta}=\bm{\beta}_{1}-\bm{\beta}_{0}\in\mathbb{R}^{p}$, the maximum variability of the response by $\Psi=\max(|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{2},|\bm{\Sigma}^{1/2}\bm{\beta}_{1}|_{2},\sigma)$, and the minimal segment length by $\Delta=\min(\theta,n-\theta)$ As a convention, we write $\theta=n$, when $\bm{\beta}_{0}=\bm{\beta}_{1}\in\mathbb{R}^{p}$, corresponding to the case of no change. By focusing on the AMOC situation under (1), we separate the two “orthogonal” aspects of the problem: The design of an adaptively optimal procedure for change point detection and estimation, which is the main focus of this paper, and the development of algorithms for multiple change point detection. While our theoretical analysis is framed under independence and Gaussianity, we also discuss extensions to accommodate temporal dependence or non-Gaussianity.
We summarise our main contributions to the growing literature on change point analysis in high-dimensional linear models, which also serve to outline the structure of the paper.
@@
    Phase transition and a new notion of sparsity. We derive minimax lower bounds for both detection and estimation of the change point, as summarised in Table 1. The (nearly) matching upper and lower bounds presented therein reveal a phase transition governed by $\mathfrak{s}:=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$, which we term the inherent sparsity, rather than the standard notion of sparsity $\mathfrak{s}_{\delta}:=|\bm{\delta}|_{0}$, where $|\cdot|_{0}$ denotes the $\ell_{0}$-norm. Notably, the results also indicate that in the dense regime, where $\mathfrak{s}\geq\sqrt{p}$, consistent estimation is possible only when $p\lesssim n^{2}$. See Section 2.
    Table 1: Minimax lower bounds for the detection and estimation of a single change point under model (1), and the corresponding upper bounds and the method that attains each upper bound. Here $\sigma_{*}$ is the smallest non-zero eigenvalue of the covariance $\bm{\Sigma}$ of regressors, and $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$ is the inherent sparsity. If $\bm{\Sigma}$ is known, we can drop the dependence on $\sigma_{*}$. In deriving the minimax lower bounds, we assume $\min\{\mathfrak{s}^{2}\log^{2}(p),\,p\}\lesssim\Delta$ and $\text{rank}(\bm{\Sigma})\asymp p$.
    Detection requisite on $\Psi^{-2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta$ Estimation error $|\widehat{\theta}-\theta|$
    Regime Lower bound Upper bound Lower bound Upper bound Method
    $\mathfrak{s}<\sqrt{p}$
    ${\mathfrak{s}}\log(\tfrac{ep}{\mathfrak{s}^{2}})$
    $\sigma_{*}^{-1}\mathfrak{s}\log(p)$
    $\frac{{\Psi^{2}}}{|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\mathfrak{s}\log(\tfrac{ep}{\mathfrak{s}^{2}})$ $\frac{\Psi^{2}}{\sigma_{*}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\mathfrak{s}\log(p)$
    McScan
    (Section 3.1)
    $\mathfrak{s}\geq\sqrt{p}$
    $\sqrt{p}$
    $\sigma_{*}^{-1}\sqrt{p\log\log(n)}$
    $\frac{{\Psi^{2}}}{|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\sqrt{p}$ $\frac{\Psi^{2}}{\sigma_{*}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}}\cdot\sqrt{p\log\log(n)}$
    QcScan
    (Section 3.2)
@@
    The first method for consistent change point detection in the dense regime. We first show that McScan (Cho et al.,, 2025), which scans the $\ell_{\infty}$-aggregation of the contrasts in local sample covariance between regressors and responses, is minimax optimal in the sparse regime up to the factor of $\sigma_{*}^{-1}$ (Section 3.1), where $\sigma_{*}$ denotes the smallest non-zero eigenvalue of $\bm{\Sigma}$, i.e. McScan achieves optimality in the sparse regime even when $\bm{\Sigma}$ is rank deficient. For the dense regime, we introduce the quadratic covariance scanning (QcScan) that evaluates the $\ell_{2}$-aggregation of the local sample covariance contrasts. QcScan is the first method that achieves minimax optimality, up to a factor of $\sigma_{*}^{-1}\sqrt{\log\log(n)}$, in detection and estimation of the change point in the dense regime even when $\bm{\Sigma}$ is rank deficient, which is in contrast to existing change point methods that require $\mathsf{Cov}(\mathbf{x}_{t})$ to have bounded spectrum. As a by-product, we also obtain an estimator for $|\bm{\Sigma}\bm{\delta}|_{2}^{2}$, which serves as a surrogate of the strength of the signal $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$ for the change point problem in (1). See Section 3.2.
@@
For a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, we denote by $\|\mathbf{A}\|$ and $\|\mathbf{A}\|_{F}$ its operator and Frobenius norms, respectively. For $0\leq a<b\leq m$, we write $\mathbf{A}_{a,b}$ to denote a sub-matrix of $\mathbf{A}$ containing its $i$th rows for $a+1\leq i\leq b$; when $\mathbf{A}$ is a vector, $\mathbf{A}_{a,b}$ denotes the corresponding sub-vector. For sequences $\{a_{m}\}$ and $\{b_{m}\}$ of positive numbers, we write $a_{m}\lesssim b_{m}$ or equivalently $a_{m}=O(b_{m})$, if $a_{m}\leq Cb_{m}$ for some finite constant $C>0$. If $a_{m}\lesssim b_{m}$ and $b_{m}\lesssim a_{m}$, we write $a_{m}\asymp b_{m}$. We write $a\vee b=\max(a,b)$ for $a,b\in\mathbb{R}$. By $\mathbf{0}$ and $\mathbf{I}$, we denote the vector of zeros and the identity matrix, respectively, whose dimensions depend on the context.
2 Minimax lower bounds
@@ 2.1 Isotropic case
Let us define a model space under (1) as
$\mathcal{P}_{\bm{\Sigma},\sigma}^{\mathfrak{s},n,p}(\tau)=\Bigl{\{}\mathsf{P}_{\theta,\bm{\beta}_{0},\bm{\beta}_{1}}\;:\;\{(\mathbf{x}_{t},Y_{t})\}_{t\in[n]}\sim\mathsf{P}_{\theta,\bm{\beta}_{0},\bm{\beta}_{1}}\text{ \ such that \ }\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})\text{ and}\\ \text{independently \ }\varepsilon_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2}),\;Y_{t}=\mathbf{x}_{t}^{\top}\bigl{(}\bm{\beta}_{0}\cdot\mathbb{I}_{\{t\leq\theta\}}+\bm{\beta}_{1}\cdot\mathbb{I}_{\{t>\theta\}}\bigr{)}+\varepsilon_{t},\\ \bm{\delta}=\bm{\beta}_{1}-\bm{\beta}_{0},\,\Delta=\min(\theta,n-\theta),\,\,\Psi=\max(|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{2},|\bm{\Sigma}^{1/2}\bm{\beta}_{1}|_{2},\sigma),\\ |\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=\mathfrak{s}\text{ \ and \ }|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta\geq\tau\Psi^{2}\mathfrak{s}\log\bigl{(}1+\tfrac{p}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p}}{\mathfrak{s}}\bigr{)}\Bigr{\}}.$
@@
Let $\sigma>0$ and $1\leq\Delta\leq n/4$, and suppose that $\bm{\Sigma}=\mathbf{I}$.
  1. (i)
    If $\mathfrak{s}\log\bigl{(}1+\tfrac{p}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p}}{\mathfrak{s}}\bigr{)}\leq\sqrt{\tfrac{25}{2}\log\left(\tfrac{17}{12\sqrt{2}}\right){\Delta}}$ and $0<\tau\leq\tfrac{2}{5}\log\left(\tfrac{17}{12\sqrt{2}}\right)$, then
@@
    $\inf_{\widehat{\theta}}\;\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{\mathfrak{s},n,p}(\tau)}\;\mathsf{P}\left(\bigl{|}\widehat{\theta}-\theta(\mathsf{P})\bigr{|}\geq{\Delta}\right)\;\geq\;\frac{1}{25}.$
  2. (ii)
    Let $\tau$ be a large enough constant, or $\tau\to\infty$ (at an arbitrary rate) as $n\to\infty$. If $\mathfrak{s}\log\bigl{(}1+\tfrac{p}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p}}{\mathfrak{s}}\bigr{)}\leq\sqrt{\tfrac{5}{2}\log\left(\tfrac{3}{2\sqrt{2}}\right)\tfrac{\Delta}{\tau}}$, then
    $\inf_{\widehat{\theta}}\;\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{\mathfrak{s},n,p}(\tau)}\;\mathsf{P}\left(\bigl{|}\widehat{\theta}-\theta(\mathsf{P})\bigr{|}\geq\frac{\sigma^{2}\mathfrak{s}\log\left(\frac{3}{2\sqrt{2}}\right)\log\left(1+\frac{p}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p}}{\mathfrak{s}}\right)}{5|\bm{\delta}|_{2}^{2}}\right)\;\geq\;\frac{1}{4}.$
Theorem 1 provides the minimax lower bounds on both the detection boundary and the rate of estimation for the change point problem in (1), which reveals a phase transition with respect to the sparsity. In an isotropic Gaussian setting with $\bm{\Sigma}=\mathbf{I}$, the sparsity is determined by $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=|\bm{\delta}|_{0}=\mathfrak{s}_{\delta}$, and the minimax lower bounds undergo a phase transition at $\mathfrak{s}=\sqrt{p}$, from that $\mathfrak{s}\log\bigl{(}1+\tfrac{p}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p}}{\mathfrak{s}}\bigr{)}\asymp\mathfrak{s}\log(ep/\mathfrak{s}^{2})\mathbb{I}_{\{\mathfrak{s}<\sqrt{p}\}}+\sqrt{p}\;\mathbb{I}_{\{\mathfrak{s}\geq\sqrt{p}\}}$. This phenomenon has hitherto been unknown in the literature, with the existing minimax lower bounds focusing on the sparse regime of $\mathfrak{s}<\sqrt{p}$ (Rinaldo et al.,, 2021; Cho et al.,, 2025). Notably, one striking observation in the dense regime is that the minimax rate of estimation becomes a trivial one when $p\gtrsim n^{2}$. This stands in sharp contrast to the sparse regime, where consistent change point detection remains feasible even in ultra-high dimensions.
@@
Figure 1 numerically illustrates the phase transition as the sparsity ($x$-axis, in a log scale) and the sample size vary (left to right), while $|\bm{\delta}|_{2}=8$, $p=900$ and other model parameters remains constant and $\bm{\Sigma}=\mathbf{I}$. The figure reports the scaled estimation errors ($y$-axis, in a square root scale) attained by various change point detection methods, including McScan (Section 3.1; a variant of Cho et al.,, 2025) and QcScan (Section 3.2) which are shown to be minimax (near-)optimal in the sparse and the dense regimes, respectively. While QcScan’s estimation error remains roughly constant regardless $\mathfrak{s}$, McScan achieves a smaller error for the sparse situations which increases with $\mathfrak{s}$ until it cross that of QcScan near $\mathfrak{s}=\sqrt{p}$; similar behaviour is also observed from the competing methods (MOSEG and CHARCOAL) tailored for the sparse regime. The OcScan (Section 4.1) which combines the outputs from McScan and QcScan, and its refinement OcScan.R (Section 5), are able to adaptively enjoy the optimal performance of McScan and QcScan regardless of the regimes determined by $\mathfrak{s}$. In particular, OcScan.R improves upon the estimators of McScan and OcScan in the sparse regime, which demonstrates the efficacy of the refinement proposed in Section 4.2.

### v3

@@ Abstract
This paper investigates the detection and estimation of a single change in high-dimensional linear models. We derive the minimax lower bound for the detection boundary which uncovers a phase transition governed by the sparsity of the covariance-weighted differential parameter. This form of inherent sparsity captures a delicate interplay between the covariance structure of the regressors and the change in regression coefficients on the detectability of a change point. Complementing the lower bound, we introduce two covariance scanning-based methods, McScan and QcSan, which achieve minimax optimal detection performance (up to possible logarithmic factors) in the sparse and the dense regimes, respectively. In particular, QcScan is the first method to achieve consistency in the dense regime and further, we devise a combined procedure which is adaptively minimax optimal across sparse and dense regimes without the knowledge of the sparsity. Computationally, covariance scanning-based methods avoid costly computation of Lasso-type estimators and attain worst-case computation complexity that is linear in the dimension and sample size. Additionally, we consider the post-detection estimation of the differential parameter and the refinement of the change point estimator. Simulation studies support the theoretical findings and demonstrate the computational and statistical efficiency of the proposed covariance scanning methods.
@@ 1 Introduction
The change point problem in linear models has received great attention over the past several decades, owing to its widespread applicability across diverse scientific and practical disciplines, including genomics, climatology, finance and economics. Numerous methods have been developed that achieve consistency in (multiple) change point detection, both in multivariate settings (e.g. Csörgo and Horváth, 1997; Bai and Perron, 1998) and high-dimensional scenarios (e.g. Lee et al., 2016; Kaul et al., 2019; Wang et al., 2021b; Rinaldo et al., 2021; Cho and Owens, 2024; Xu et al., 2024).
When faced with high dimensionality, many existing methods rely on $\ell_{1}$-regularised estimators of local regression parameters for change point detection, thereby requiring the sparsity of regression parameters. Gao and Wang (2022) propose a projection-based method that instead assumes sparsity on the differences of regression parameters. However, it is applicable only when the dimension is strictly smaller than the sample size. More recently, Cho et al. (2025) introduce McScan, a multiscale covariance scanning approach that achieves consistent change point detection with both statistical and computational efficiency. McScan does not require exact sparsity of the regression parameters or their differences, and accommodates ultra high-dimensional settings. It operates by examining the $\ell_{\infty}$-norm of the contrasts in local sample covariances between the response and the regressors, and thus is better suited to sparse regimes implicitly. We also remark that minimax lower bounds for the change point problem have been investigated, but only when the sparsity of the regression parameters (or their differences) is bounded by the square root of the dimension (see Rinaldo et al., 2021 and Cho et al., 2025).
Despite these advances, several questions remain to be answered for the change point problem in high-dimensional linear models. In particular, it is hitherto unknown whether it is possible to consistently detect the change point when the change is non-sparse. More fundamentally, it is yet to be understood which notion of sparsity — or more generally, which structural assumption — determines the intrinsic difficulty of the change point problem. We provide first answers to these questions under a Gaussian linear model with at most one change (AMOC): The observations $(Y_{t},\mathbf{x}_{t})\in\mathbb{R}^{1+p},\,t\in[n]:=\{1,\ldots,n\}$, are distributed as
@@
where $\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})$ with non-negative definite $\bm{\Sigma}\in\mathbb{R}^{p\times p}$ and independently, $\varepsilon_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2})$ for all $t$. We denote by $q\in\{0,1\}$ the number of change points. In the case of a single change point ($q=1$), we define the differential parameter as $\bm{\delta}=\bm{\beta}_{1}-\bm{\beta}_{0}\in\mathbb{R}^{p}$, the maximum variability of the response as $\Psi=\max(|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{2},|\bm{\Sigma}^{1/2}\bm{\beta}_{1}|_{2},\sigma)$, and the minimal segment length as $\Delta=\min(\theta,n-\theta)$. In the absence of any change point ($q=0$), where $\bm{\beta}_{0}=\bm{\beta}_{1}$, we write $\theta=n$ as a convention.
Under the model (1), we are interested in the problem of consistently detecting the presence of any change point and, if $q=1$, estimating its location by a consistent estimator $\widehat{\theta}$ satisfying $|\widehat{\theta}-\theta|/\Delta\to 0$ with probability tending to one as $n\to\infty$. By focusing on the AMOC situation, we separate the two “orthogonal” aspects of the problem: The design of an adaptively optimal procedure for change point detection, which is the main focus of this paper, and the development of algorithms for multiple change point detection. While our theoretical analysis is framed under independence and Gaussianity, we also discuss extensions to accommodate temporal dependence or non-Gaussianity.
We summarise our main contributions to the growing literature on change point analysis in high-dimensional linear models, which also serves to outline the structure of the paper.
@@
    Phase transition and a new notion of sparsity. We derive the minimax lower bounds for the change point detection as summarised in Table 1. With the (nearly) matching upper bounds presented therein, it reveals a phase transition governed by $\mathfrak{s}:=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$, which we term the inherent sparsity, rather than the standard notion of sparsity $\mathfrak{s}_{\delta}:=|\bm{\delta}|_{0}$, where $|\cdot|_{0}$ denotes the $\ell_{0}$-semi-norm. See Section 2.
    Table 1: Minimax lower bound on $\Psi^{-2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta$, for the detection of a single change point under the model (1), corresponding upper bounds and the methods that attain the upper bounds. Here $\sigma_{*}$ is the smallest non-zero eigenvalue of the covariance $\bm{\Sigma}$ of regressors, $\sigma_{X}^{2}$ is the largest diagonal entry of $\bm{\Sigma}$, and $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}$ is the inherent sparsity. If $\bm{\Sigma}$ is known, we can drop the dependence on $\sigma_{*}$, $\sigma_{X}^{2}$ and $\|\bm{\Sigma}\|$. In deriving the lower bound, we assume $\min\{\mathfrak{s}^{2}\log^{2}(p),\,p\}\lesssim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n^{\nu}}$, $\nu\in(0,1/2)$, and $\text{rank}(\bm{\Sigma})\asymp p$.
    Regime Lower bound (1) Upper bound (Theorems 3 and 5) Method
    $\mathfrak{s}<\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$ ${\mathfrak{s}}\log(\tfrac{ep{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}{\mathfrak{s}^{2}})$ $\tfrac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}}{\sigma_{*}}\mathfrak{s}\log(p\log(n))$ McScan (Section 3.1)
    $\mathfrak{s}\geq\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log n}}$ $\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$ $\tfrac{\|\bm{\Sigma}\|}{\sigma_{*}}\sqrt{p\log\log(n)}$ QcScan (Section 3.2)
@@
    The first method for consistent change point detection when $\bm{\Sigma}$ is rank deficient and in the dense regime. We first show that McScan (Cho et al., 2025), which scans the $\ell_{\infty}$-aggregation of the contrasts in local sample covariance between the response and the regressors, is minimax optimal in the sparse regime up to the factor of $\sigma_{*}^{-1}\sigma_{X}^{2}$ and possible logarithmic factors in $p$ and $n$ (Section 3.1), where $\sigma_{*}$ denotes the smallest non-zero eigenvalue of $\bm{\Sigma}$ and $\sigma_{X}^{2}$ its largest diagonal entry. Additionally, we introduce the quadratic covariance scanning (QcScan) that evaluates the $\ell_{2}$-aggregation of the local sample covariance contrasts, and show QcScan to be the first method that achieves minimax optimality, up to a factor of $\sigma_{*}^{-1}\|\bm{\Sigma}\|$, in the dense regime. In short, both McScan and QcScan accommodate rank-deficient $\bm{\Sigma}$, which is in contrast to existing change point methods that require $\mathsf{Cov}(\mathbf{x}_{t})$ to have bounded spectrum. As a by-product, we also obtain an estimator for $|\bm{\Sigma}\bm{\delta}|_{2}^{2}$, which serves as a surrogate of the strength of the signal $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$ for the change point problem in (1). See Section 3.2.
@@
For a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, we denote by $\|\mathbf{A}\|$ and $\|\mathbf{A}\|_{F}$ its operator and Frobenius norms, respectively. For $0\leq a<b\leq m$, we write $\mathbf{A}_{a,b}$ to denote a sub-matrix of $\mathbf{A}$ containing its $i$th rows for $a+1\leq i\leq b$; when $\mathbf{A}$ is a vector, $\mathbf{A}_{a,b}$ denotes the corresponding sub-vector. By $\lvert\cdot\rvert_{\nu}$ we denote the $\ell_{\nu}$-(semi-)norm of a vector for $\nu\in[0,\infty]$. For sequences $\{a_{m}\}$ and $\{b_{m}\}$ of positive numbers, we write $a_{m}\lesssim b_{m}$ or equivalently $a_{m}=O(b_{m})$, if $a_{m}\leq Cb_{m}$ for some finite constant $C>0$. If $a_{m}\lesssim b_{m}$ and $b_{m}\lesssim a_{m}$, we write $a_{m}\asymp b_{m}$. We write $a\vee b=\max(a,b)$ for $a,b\in\mathbb{R}$. By $\mathbf{0}$ and $\mathbf{I}$, we denote the vector of zeros and the identity matrix, respectively, whose dimensions depend on the context.
2 Minimax lower bound
@@ 2.1 Isotropic case
Let us define a model space with a single change under (1), as
$\mathcal{P}_{\bm{\Sigma},\sigma}^{\mathfrak{s},n,p}(\tau)=\biggl\{\mathsf{P}_{\theta,\bm{\beta}_{0},\bm{\beta}_{1}}\;:\;\{(Y_{t},\mathbf{x}_{t})\}_{t\in[n]}\sim\mathsf{P}_{\theta,\bm{\beta}_{0},\bm{\beta}_{1}}\text{ \ such that \ }\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})\text{ and}\\ \text{independently \ }\varepsilon_{t}\sim_{\text{iid}}\mathcal{N}(0,\sigma^{2}),\;Y_{t}=\mathbf{x}_{t}^{\top}\bigl(\bm{\beta}_{0}\cdot\mathbb{I}_{\{t\leq\theta\}}+\bm{\beta}_{1}\cdot\mathbb{I}_{\{t>\theta\}}\bigr)+\varepsilon_{t},\\ \bm{\delta}=\bm{\beta}_{1}-\bm{\beta}_{0},\,\Delta=\min(\theta,n-\theta),\,\,\Psi=\max(|\bm{\Sigma}^{1/2}\bm{\beta}_{0}|_{2},|\bm{\Sigma}^{1/2}\bm{\beta}_{1}|_{2},\sigma),\\ |\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=\mathfrak{s}\text{ \ and \ }|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta\geq\tau\Psi^{2}\mathfrak{s}\log\Bigl(1+\tfrac{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}}{\mathfrak{s}}\Bigr)\biggr\},$
with some constant $\tau>0$. We also define the model space without any change point as
[…]
@@
Let $\sigma>0$ and $n$ be sufficiently large. We consider the case where $\bm{\Sigma}=\mathbf{I}$. If
$\mathfrak{s}\log\left(1+\frac{p\log\log(n)}{\mathfrak{s}^{2}}\vee\frac{\sqrt{p\log\log(n)}}{\mathfrak{s}}\right)\leq n^{\nu}$
for some $\nu\in(0,1/2)$, then it holds, for every $\tau\in(0,1/6]$,
@@
$\inf_{\psi}\biggl\{\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{0,n,p}}\mathsf{E}_{\mathsf{P}}\left[\psi\left((Y_{t},\mathbf{x}_{t})_{t\in[n]}\right)\right]+\sup_{\mathsf{P}\in\mathcal{P}_{\mathbf{I},\sigma}^{\mathfrak{s},n,p}(\tau)}\mathsf{E}_{\mathsf{P}}\left[1-\psi\left((Y_{t},\mathbf{x}_{t})_{t\in[n]}\right)\right]\biggr\}\geq\frac{1}{4e},$
where the infimum is taken over all measurable functions $\psi:\mathbb{R}^{n\times(1+p)}\to[0,1]$.
Theorem 1 provides the minimax lower bound on the detection boundary for the change point problem in (1), which reveals a phase transition with respect to the sparsity. In an isotropic Gaussian setting with $\bm{\Sigma}=\mathbf{I}$, the sparsity is determined by $\mathfrak{s}=|\bm{\Sigma}^{1/2}\bm{\delta}|_{0}=|\bm{\delta}|_{0}=\mathfrak{s}_{\delta}$, and the minimax lower bound undergoes a phase transition at $\mathfrak{s}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$, from that
$\mathfrak{s}\log\left(1+\tfrac{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}{\mathfrak{s}^{2}}\vee\tfrac{\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}}{\mathfrak{s}}\right)\asymp\begin{cases}\mathfrak{s}\log(ep{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}/\mathfrak{s}^{2})&\text{ if }{\mathfrak{s}<\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}},\\ \sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}&\text{ if }{\mathfrak{s}\geq\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}}.\end{cases}$
This phenomenon has hitherto been unknown in the literature, with the existing investigation into minimax lower bounds focusing on the sparse regime of $\mathfrak{s}<\sqrt{p}$ (Rinaldo et al., 2021; Cho et al., 2025).
@@
Figure 1 numerically illustrates the phase transition as the sparsity ($x$-axis, in a log scale) and the sample size vary (left to right) in the isotropic setting. The figure reports the scaled estimation errors ($y$-axis, in a square root scale) attained by various change point detection methods, including McScan (Section 3.1; a variant of Cho et al., 2025) and QcScan (Section 3.2) which are later shown to be minimax (near-)optimal in the sparse and the dense regimes, respectively. While QcScan’s estimation error remains roughly constant regardless $\mathfrak{s}$, McScan attains a smaller error for the sparse situations which increases with $\mathfrak{s}$ until it crosses that of QcScan near $\mathfrak{s}=\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$; similar behaviour is also observed from the competing methods (MOSEG and CHARCOAL) tailored for the sparse regime. The OcScan (Section 4.1) which combines the outputs from McScan and QcScan, and its refinement OcScan.R (Section 5), are able to adaptively enjoy the optimal performance of McScan and QcScan regardless of the regimes determined by $\mathfrak{s}$. In particular, OcScan.R improves upon the estimators of McScan and OcScan in the sparse regime, which demonstrates the efficacy of the refinement proposed in Section 4.2.